OpenStudy (anonymous):

If `f(x) = (x +1)^-1` and `g(x) = x - 2`, what is the domain of `f(x)-: g(x)`? all values of x `(oo, -1), (-1,2)`, and `(2, oo)` `(-oo,2)` and `(2,oo)` `(-oo,-1]` and `[2,oo)`

hero (hero):

I assume you were trying to create this: f(x) ÷ g(x)

OpenStudy (anonymous):

yes, my computer is messing up a little

hero (hero):

alt 0247 to create ÷

hero (hero):

So basically: \[\f(x) \div g(x) = \frac{1}{x+1} \div (x - 2) = \frac{1}{x+1} \dot\ \frac{1}{x-2} = \frac{1}{(x+1)(x-2)}\]

hero (hero):

Finish the rest